City planners wish to estimate the mean lifetime of the most commonly planted trees in urban settings. A sample of 16 recently felled trees yielded mean age 32.7 years with standard deviation 3.1 years. Assuming the lifetimes of all such trees are normally distributed, construct a 99.8% confidence interval for the mean lifetime of all such trees.


1
Expert's answer
2022-06-03T13:22:04-0400

The critical value for "\alpha = 0.002, df=n-1=15" degrees of freedom is "t_c\u200b=z_{1\u2212\u03b1\/2;n\u22121}= 3.73283"

The corresponding confidence interval is computed as shown below:



"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(32.7- 3.73283\times\dfrac{3.1}{\sqrt{16}},""32.7+ 3.73283\times\dfrac{3.1}{\sqrt{16}})"




"=(29.807, 35.593)"

Therefore, based on the data provided, the 99.8% confidence interval for the population mean is "29.807 < \mu < 35.593," which indicates that we are 99.8% confident that the true population mean "\mu" is contained by the interval "(29.807, 35.593)."